Given the following total revenue (TR) and total cost(TC) equations, determine the output rate that would result in a breakeven(i.e.,zero profit) situation for a firm
TR = 51Q - Q^2
TC = 625 + Q
1
Expert's answer
2018-03-04T09:47:07-0500
Breakeven situation is when total revenue is similar to the total cost
TR = TC
51Q - Q^2 = 625 + Q
51Q - Q^2 – 625 – Q = 0
-Q^2+50Q – 625 = 0
Q^2 – 50Q + 625 = 0
Q = 25